# Superconducting qubits at the utility scale: the potential and limitations of modularity

### v2

@@ I.4.2 Stitching strategy
Following the widgetization process, we can construct an approximate version of $S_{\text{est}}$. We propose the widgetization-based set $\tilde{S}_{\text{est}}$ as a good approximation of $S_{\text{est}}$ for resource estimation purposes, defined as follows:
@@
Recall that $n^{T}_{\text{init}}$ and $n^{R_{Z}}_{\text{init}}$ are the exact number of explicit $\{T,T^{\dagger}\}$ and arbitrary-angle $R_{Z}$ operations in $U$. In this widgetization-based estimation procedure, we first obtain estimation primitives for each widget based on its graph, sometimes referred to as a subgraph. We can find the overall circuit resource estimates as follows:
@@
Equation 10 emphasizes the key assumptions we made to generate our estimates, formally outlining our stitching strategy. I.e.  how we combine the individual graphs and scheduling lists to approximate the complete $\mathcal{G}$, as well as $S^{\mathcal{G}}_{\text{prep}}$, $S^{\mathcal{G}}_{\text{consump}}$, and $S^{\mathcal{G}}_{\text{meas}}$. While an FTQC needs to physically stitch the widgets (through teleportations) to execute the complete circuit, the same is unnecessary for resource estimation purposes. Consequently, we reduced the computational demands for obtaining resource estimates by calculating the widgetized quantities in eq. 10 through software.
I.4.3 Pre-processing the input subcircuits
To prepare for graph compilation, we must parse each widget and transpile it to include only logical Clifford+$T$+$R_{Z}(\theta)$ gates, using exact gate equivalences. Here, the $\theta$ values represent non-Clifford angles. Later, during the consumption stage, when we want to perform single-qubit logical measurements, we need to further decompose each $R_{Z}(\theta)$ node into a sequence of Clifford+$T$ gates through gate synthesis [68, 69].
@@ I.4.3 Pre-processing the input subcircuits
Once transpilation is complete, the next step is to generate the unified MBQC-based set of graph states $\{g\}$ and schedules $\{S^{g}_{\text{prep}}\}$, $\{S^{g}_{\text{consump}}\}$, and $\{S^{g}_{\text{meas}}\}$ (and other essential intermediary lists detailed below), which leads to $\tilde{S}_{\text{est}}$ sufficient for the resource estimations of $U$.
I.4.4 Cabaliser: compiling to graphs and consumption schedules
RRE utilizes the stabilizer tableau-based fault-tolerant compiler Cabaliser [71] as a crucial intermediate layer to transform the set of individual logical widgets, $\{U\}$, into corresponding MBQC-based subgraphs, consumption schedules, “locally simulated” quantum states, and other essential information. Locally simulated quantum states are Cabaliser’s outputs derived through exact calculations. Cabaliser effectively maps an input state |$\ket{\text{state}}_{i-2}$ into |$\ket{\text{state}}^{\text{Cabaliser}}_{i}$: it applies $U_{i}$, Pauli corrections, the consumption subschedule, and additional logical operations (such as Hadamards) to ($i-1$)th graph state to create a faithful vector representation of the corresponding universal subgraph at step $i$. It is guaranteed that |$\ket{\text{output}}\equiv\ket{\text{state}}^{\text{Cabaliser}}_{n_{\text{widget}}-1}$. This vector representation helps verify the unitaritiness of RRE and Cabaliser’s outputs for small $n_{\text{input}}$. We present a systematic method to verify the graphs and schedules produced by RRE, taking into account widgetization and stitching in section A.3. For complete details on scalability and various compilation components of Cabaliser, we refer readers to [71, 38, 31].
For a given individual subcircuit $U_{i}$, which includes only Clifford+$T$+$R_{Z}(\theta)$, Cabaliser can generate the following set of quantities:
@@ I.4.4 Cabaliser: compiling to graphs and consumption schedules
In this equation, $S_{\text{frames}}^{g_{i}}$ is a set that tracks the required Pauli corrections (frames) to be applied to the measurements at the end. $S_{\text{output}}^{g_{i}}$ and $S_{\text{input}}^{g_{i}}$ represent the sets (or maps) of output and input nodes in the graph state, respectively. Since we decompose the entire algorithm in the time direction with a uniform width that covers all qubits, the number of intermediate input and output nodes remains the same. $n^{g_{i}}_{\text{logical}}$ denotes the quantum memory needed to process $U_{i}$. $S_{\text{consump}}^{g_{i}}$ identifies the graph consumption schedule for performing individual logical qubit measurements. $S_{\text{meas}}^{g_{i}}$ identifies which rotational basis we must use for each measurement during the consumption stage. Finally, |$\ket{\text{state}}^{\text{Cabaliser}}_{i}$ represents the locally simulated quantum state extracted from the outputs generated up to $i$th step.
We repeat the process of eq. 11 for all $n^{*}_{\text{widgets}}$ distinct widgets, providing the user with the information required for the approximate estimation set in eq. 9, excluding the preparation schedules, $\{S^{g}_{\text{prep}}\}$, which we generate using another tool detailed below.
I.4.5 Substrate Scheduler: generating preparation schedules
RRE utilizes the Substrate Scheduler [44] to map the graph states $\{g\}$, calculated by Cabiliser in eq. 11, to a set of graph preparation sub-schedules $\{S_{\text{prep}}^{g}\}$. These preparation sub-schedules are optimized for the quantum memory register and ancilla bus in the superconducting hardware layout described in fig. 1.
For the Substrate Scheduler calculations, we assume a dual-rail or bilinear quantum memory and ancilla bus pattern consisting of linear arrangements of surface code qubits for storing graphs and mediating logical measurements on the nodes as per fig. 1. The pattern has a fixed size $\tilde{n}_{\text{logical}}:=\max_{g_{i}}(n^{g_{i}}_{\text{logical}})$ per rail. This assumption ensures no memory constraints when executing the preparation and consumption stages consecutively. While the intra-module architecture we use for resource estimates follows fig. 6, the Substrate Scheduler results can still be matched and utilized for the wrapped bilinear pattern. Currently, the Substrate Scheduler identifies an optimal schedule based on the individual mapping of each $g$ to this layout. The Substrate Scheduler’s performance could be enhanced by incorporating additional information about node types and supporting customized bus architectures, which we plan to address in future work.
I.4.6 Space-time-optimal resource estimation
DistillWidget (superconducting) $p_{out}$ Phys. Space($L_{\text{width}}\times L_{\text{length}}$) Qubits, $Q~{}$ Cycles, $C$
(15-to-1)17,7,7 $4.5\times 10^{-8}$ 64$\times$72 4620 42.6
@@ I.4.6 Space-time-optimal resource estimation
Table 1: Distillation Widget Lookup Table: Litinski’s distillation widgets that match our superconducting architecture with $p=10^{-3}$. Extended from [11, Table 1].
@@
At a macroscopic level, the architecture features a ladder-style structure with $n_{\text{per-leg}}^{\text{modules}}$ pairs of QPU modules connecting the legs, as illustrated in fig. 3. We based this design on the assumption that superconducting platforms have a low syndrome extraction cycle and would benefit from minimizing the physical footprint while interleaving graph preparation and consumption. Consequently, we always create $\tilde{n}_{\text{logical}}$ surface code patches for logical memory patches shared among all modules on a leg of the macro-architecture.
At the intra-module level, we organize each module as a square grid of logical qubits (patches), the foundational building blocks for all components. The edge size is determined by $l_{\text{edge}}=\lfloor\sqrt{\frac{N_{\text{phys}}}{2d^{2}}}\rfloor$, where $N_{\text{phys}}=1,000,000$ represents the total number of physical qubits available per module. We assumed each rotated surface code patch contains $2d^{2}$ physical qubits.
We performed a static allocation of intra-modular components, providing sufficient resources for the largest widgets as illustrated in fig. 6. Each module has five main components to which we allocate physical qubits. The unallocated or "wasted" space can still be utilized for some ancillary operations. The first component is a quantum register comprising memory or graph-states qubits. The second is a chain of ancilla bus qubits of the same length, facilitating logical measurements on the memory qubits, which connects to the third component, i.e., the supplementary bus. The first two components are arranged in a snake-like pattern to occupy one QPU edge fully. The fourth component is the linear $T$-transfer bus to supply the ancilla bus portion with $T$-states for graph consumption. The last is $T$-distilleries, also called distillation widgets or factories. See section I.3.2 for details.
The space-time volume of the complete graph creation on the bi-linear quantum bus before single-qubit logical measurements is
@@
[…]
In “Distillation Widget Lookup Table”, table 1, the output probability, denoted as $p_{\text{out}}$, determines the error rate for the ancillary state of factories and sets an upper limit on the error rate for any $T$-gate implemented using the ancilla. The specifications for the number of qubits, the dimensions of the 2D patch, and the number of physical cycle sweeps all depend on the construction of the factory. Our analysis focused exclusively on selecting distillation widgets that produce $T$-states, thus excluding any entries related to $CCZ$ from the original [11, Table 1].
Next, we must include the resource cost for decomposing non-Clifford into Clifford+$T$ (gate synthesis) to consume graph-state nodes through single-qubit measurements. After completing a graph preparation sub-schedule, at each sub-step of the consumption schedule, we utilize one or more purified $T$-resources to perform measurements on the desired basis. The FTQC must teleport the $T$-states from distillation factories to the ancilla bus. We assume that the graph nodes are either of the type that requires arbitrary $R_{Z}$-basis measurements or a single $T$-basis measurement. Measurements that involve $R_{Z}$-axis rotations are decomposed into sequences of $T$-gates, with a fixed worst-case length of $L_{\varepsilon}=\lceil c_{0}\log(1/\varepsilon)+c_{1}\rceil$. We determine this length through the gate synthesis methods available to RRE. Here, $\varepsilon$ refers to the diamond-norm error for decompositions into Clifford+$T$—see [68, 69] for further details.
The generation of each distilled $T$-state in the factories requires $C$ cycles, and we need one further surface code tock (equivalent to $d$ cycles) to interact it with the graph node and measure it in either the $X$ or $Z$ basis. In this work, we ignore the time cost of any operation that requires $O(1)$ cycles (e.g., shuffling surface code patches), assuming we can perform them simultaneously during sub-steps at no additional cost. Therefore, for each non-Clifford measurement of memory nodes, assuming an $R_{Z}$-basis as an example, the volume increases as follows:
@@
In eq. 13, the second term reflects the fact that both graph node consumption and distillation occur over a space of $2\tilde{n}_{\text{logical}}+n_{\text{per-leg}}^{\text{modules}}l_{\text{transfer-bus}}$. All nodes with distinct measurement basis types, $T$ and $R_{Z}$, must be measured sequentially.
If we assume the worst-case sequential execution, the upper bound for full space-time volume becomes:
@@
We define the total number of sequential T-states for graph consumption as $N^{\text{seq}}_{\text{consump}}=\left\lceil\frac{n^{T}_{\text{init}}}{n^{\prime}}\right\rceil+L_{\varepsilon}\left\lceil\frac{n^{R_{z}}_{\text{init}}}{n^{\prime}}\right\rceil$. Recall that we assumed a simplified $T$-routing scheme that allows the ancilla bus to serve up to $n^{\prime}$ simultaneous magic states. FTQC can only consume distinct $T$ and $R_{Z}$ nodes in parallel. The number of sequential T-states to distill is $N^{\text{seq}}_{\text{distill}}=\left\lceil\frac{n^{T}_{tot}}{n^{\prime}}\right\rceil$, where the total $T$-state measurements required is defined as $n^{T}_{tot}:=n^{R_{Z}}_{\text{init}}L_{\varepsilon}+n^{T}_{\text{init}}$. The quantity $n^{T}_{tot}$ is the proxy we use for $T$-count and typically becomes larger than $N_{\mathcal{G}}$. We do not include distillation widget space-time volumes, as their errors subsume into the output probability $p_{out}$ of the distilled state [11, Fig. 1]. If the select widget from table 1 has a second (20-to-4) layer (the second and third widgets), then we must replace $n^{\prime}\rightarrow 4n^{\prime}$ in all space-time equations.
As detailed in section I.3.1, a single cycle time of the surface code, $p_{C}$, has a failure probability characterized by the physical and logical error rates relationship given in eq. 1. Consequently, we can express the failure probability of a single unit cell containing $d$ cycles as:
$p_{\text{log-cell}}=1-(1-p_{C})^{d}.$ (15)
The failure rate for a total volume of these cells, denoted as $V^{\text{sequential}}_{\text{total}}$, must satisfy the following condition:
$p_{\text{algo-fail}}>1-(1-p_{\text{log-cell}})^{V^{\text{sequential}}_{\text{total}}},$
[…]
Assuming that $p_{C}\ll 1$, we can approximate $(1-p_{C})^{d}\approx 1-p_{C}d$, leading us to conclude that $p_{\text{log-cell}}\approx p_{C}d$ and consequently,
@@
This equation highlights the relationship between the failure rates and the characteristic space-time volume for the QEC process.
Subtracting each side of section I.4.6 from 1 and taking the logarithm of both sides, we get
@@
If we substitute $V^{\text{sequential}}_{\text{total}}$ from eq. 14 into eq. 16, the primary logarithmic inequality to solve can be expressed as
@@
where $J_{1}=\log(1-p_{\text{algo-fail}})$. The minimum code distance required is the smallest integer $d$ that satisfies eq. 17. The only free variables are the cycle time of the distillation widget, $C$, and $\varepsilon$. We previously extracted $\mathbb{L}_{\text{prep}}$ and $\tilde{n}_{\text{logical}}$ from the Cabaliser and Substrate Scheduler.
@@
To determine $C$ and $\varepsilon$, we perform nested loops to iterate through widgets and then gate synthesis’ $\varepsilon$. The inner loop feeds back into the Clifford+$T$ decomposition as the failure rate of a unit cell, $p_{\text{log-cell}}$, establishes the approximation error as $\varepsilon<p_{\text{log-cell}}$. By selecting a specific widget, we fix $C$, which also determines the output error rate of the distilled state, $p_{\text{out}}$.
@@
In a compiled quantum algorithm, we aim to ensure that the distilled $T$-gates have precisely the same error rates as any other gates in the compiled circuit. A Pauli initialization or measurement occupies a single unit cell in the space-time volume. Thus, its failure rate is governed by eq. 15. This error rate must be dominant. Therefore, the error output, $p_{out}$, from the distillation circuit should be less than the failure rate of this cell (that is, we must ensure that the distillation output is at least as good as a unit cell in the space-time volume). Consequently, we first select the smallest value of $C$ from table 1 and solve eq. 17 for $d$. We then verify that, for a physical error rate of $p=10^{-3}$ and our calculated solution for $d$, whether
$p_{\text{log-cell}}=1-(1-p_{C})^{d}>p_{\text{out}}$ (18)
@@
does hold for the corresponding $p_{\text{out}}$ entry. If this condition is unmet, we must choose a better-performing $T$-distillation widget in the outer loop. We then proceed to the next entry on the list and repeat until we identify a distillation widget that performs at least as well as the failure of a logical unit cell.
@@
Once a specific distillation widget is selected, the total number of allocated physical qubits across all modules at any given time and the total intra-modular runtime (including the first subgraph preparation time at step-00, all consumption operations, and delays) can be calculated as follows:
@@
$\begin{split}N&{}_{\text{alloc-phys}}=\\ &4\tilde{n}_{\text{logical}}d^{2}+n^{\text{modules}}_{\text{per-leg}}(n_{\text{distilleries}}Q+2d^{2}l_{\text{transfer-bus}}),\\ t&{}^{\text{total}}_{\text{consump}}=\\ &8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{distill}}+t^{\text{delay}}_{\text{prep}}~{}.\end{split}$ (19)
@@
In eq. 19, $n^{\text{modules}}_{\text{per-leg}}$ refers to the number of modules per leg of the ladder macro-architecture – that is, there are $2n^{\text{modules}}_{\text{per-leg}}$ modules in total. $t$ denotes the characteristic gate time, establishing the architecture’s quantum tock for all intra-modular operations. The factor of 8 arises from the fact that in our physical architecture, each surface code sweep for syndrome extraction circuits necessitates one initialization, two Hadamards, four entangling gates, and one measurement, all assumed to take a duration of $t$ (see, for instance, [10, Fig. 9]). The temporal cost of initializing the first subgraph $g_{0}$ with $L^{0}_{\text{prep}}(S^{0}_{\text{prep}})$ is always nonzero and was explicitly added. FTQC carries out the remaining graph preparations and $T$-state distillation operations in parallel with consumption operations. There may be overall delays accumulated over different time steps, defined as $t^{\text{delay}}_{\text{distill}}=\sum_{i}t^{\text{distill-delay}}_{\text{i}}$ and $t^{\text{delay}}_{\text{prep}}=\sum_{i}t^{\text{prep-delay}}_{\text{i}}$, which we detail below.
@@
To execute each consumption sub-schedule $i$, a specific number of distilled magic states must exist in the module’s $T$-transfer bus. We must introduce a delay if we require more magic states than what was accumulated in the $T$-transfer bus during the prior preparation step, $i-1$. The number of $T$ states per fridge after a preparation step is
@@
$\begin{split}n^{T-\text{per-fridge}}_{i}=\max\big{(}\lfloor&\frac{n_{\text{distilleries}}t^{\text{prep}}_{i}}{8tC}\rfloor,l_{\text{tansfer-bus}}\big{)}\\ &\text{for}~{}i\in\{0,\dots,n_{\text{widgets}}-1\}.\end{split}$ (20)

### v3

@@ II.4.3 Stitching strategy
Following the widgetization process, we can construct an approximate version of $S_{\text{est}}$. We propose the widgetization-based set $\tilde{S}_{\text{est}}$ as a good approximation of $S_{\text{est}}$ for resource estimations, defined as follows:
@@
Recall that $n^{T}_{\text{init}}$ and $n^{R_{Z}}_{\text{init}}$ are the exact counts of explicit $\{T,T^{\dagger}\}$ and all arbitrary-angle $R_{Z}$ operations in $U$. In this widgetization-based estimation procedure, we first obtain estimation primitives for each widget based on its (sub)graph. We then obtain the overall circuit resource estimates as follows:
@@
Equation 10 formalizes our stitching strategy and outlines the key assumptions underlying our estimates. It identifies how we combine the individual graphs and scheduling lists to approximate the complete $\mathcal{G}$, as well as $S^{\mathcal{G}}_{\text{prep}}$, $S^{\mathcal{G}}_{\text{consump}}$, and $S^{\mathcal{G}}_{\text{meas}}$. While an FTQC must physically stitch the widgets (through teleportations) to execute the complete circuit, this is unnecessary for resource estimation. Consequently, we reduced the computational cost of obtaining resource estimates by computing the widgetized quantities in eq.˜10 through software.
II.4.4 Pre-processing the input subcircuits
@@
Once transpilation is complete, the next step is to generate the unified MBQC-based set of graph states $\{g\}$ and schedules $\{S^{g}_{\text{prep}}\}$, $\{S^{g}_{\text{consump}}\}$, and $\{S^{g}_{\text{meas}}\}$, and other essential intermediary lists detailed below, which yields $\tilde{S}_{\text{est}}$ sufficient for resource estimation of $U$.
II.4.5 Cabaliser: compiling to graphs and consumption schedules
[…]
We repeat the process of eq.˜11 for all $n^{*}_{\text{widgets}}$ distinct widgets, providing the user with the information required for the approximate estimation set in eq.˜9, excluding the preparation schedules, $\{S^{g}_{\text{prep}}\}$, which we generate using another tool detailed below.
II.4.6 Substrate Scheduler: generating preparation schedules
RRE uses the Substrate Scheduler [41] to map the graph states $\{g\}$, computed by Cabiliser in eq.˜11, to a set of graph preparation sub-schedules $\{S_{\text{prep}}^{g}\}$. These preparation sub-schedules are optimized for the logical memory register and auxiliary bus in the superconducting hardware layout shown in fig.˜1.
For the Substrate Scheduler calculations, we assume a dual-rail or bilinear quantum memory and auxiliary bus pattern consisting of linear arrangements of rotated surface code logical qubits for storing graphs and mediating logical measurements on the nodes, as shown in fig.˜1. The pattern has a fixed size $\tilde{n}_{\text{logical}}:=\max_{g_{i}}(n^{g_{i}}_{\text{logical}})$ per rail. This assumption ensures no memory constraints when executing the preparation and consumption stages consecutively. Although the intra-module micro-architecture used for resource estimates follows the layout shown in fig.˜1, the Substrate Scheduler results can still be matched and used under the assumption of a wrapped-bilinear pattern. Currently, the Substrate Scheduler identifies an optimal schedule based on the individual mapping of each $g$ to this layout. The Substrate Scheduler’s performance could be enhanced by incorporating additional information about node types and supporting customized bus architectures, which we plan to address in future work.
II.4.7 Space-time-optimal resource estimation
$T$-factory(superconducting) $p_{out}$ Phys. Space($L_{\text{width}}\times L_{\text{length}}$) Physical Qubits, QQ~ Cycles, $C$
@@ II.4.7 Space-time-optimal resource estimation
Table 1: $T$-factory Lookup Table: Litinski’s $T$-factories that match our superconducting architecture with $p=10^{-3}$. Extended from [40, table 1].
In this section, we formulate the space-time volume requirements for graph state processing parts based on $\{\text{config}\}$, $\tilde{S}_{\text{est}}$, and the architectures presented in section˜II.3, with a spatially optimal size of the quantum register and time-optimal compilation strategies (given the spatial configurations), creating a “middle-of-the-way" design.
At a macroscopic level, the architecture features a ladder-style structure with $n_{\text{per-leg}}^{\text{modules}}$ pairs of QPU modules connecting the legs, as illustrated in fig.˜3. We based this design on the assumption that superconducting platforms have a low syndrome extraction cycle and would benefit from minimizing the physical footprint while interleaving graph preparation and consumption. Consequently, we always create $\tilde{n}_{\text{logical}}$ rotated surface code patches for the logical memory register shared among all modules on a leg of the macro-architecture.
At the intra-module level, we organize each module as a square grid of logical qubits (patches), the foundational building blocks for all components. The edge size is determined by $l_{\text{edge}}=\left\lfloor\sqrt{\frac{n_{\text{phys}}}{2d^{2}}}\right\rfloor$, where $n_{\text{phys}}=1,000,000$ represents the total number of physical qubits available per module. For simplicity in resource estimates, we assumed each rotated surface code patch contains $2d^{2}$ physical qubits.
We performed a static allocation of intra-modular components, providing sufficient resources for the largest widgets, as illustrated in fig.˜1. Each module has five main components, to which we allocate physical qubits. The unallocated or “wasted” space can still be used for ancillary operations. The first component is a quantum register of logical memory (graph-state) qubits. The second is a chain of logical auxiliary bus qubits of the same length, facilitating logical measurements on the memory qubits, and connects to the third component, i.e., the supplementary bus. The first two components are arranged in a snake-like bilinear pattern to fully occupy one QPU edge. The fourth component is the linear $T$-transfer bus, which supplies the auxiliary bus portion with $T$-states for graph consumption. The last components are $T$-factories. See section˜II.3.2 for details.
The space-time volume of the complete graph creation on the bilinear pattern before single-qubit logical measurements is
@@
We use the same $T$-factory, matched from an extended list of appropriate widgets, across all steps and modules. We build the extended $T$-factory list using a method proposed by Litinski (see the original table 1 in [40]). We consistently assume a homogeneous superconducting physical error rate of $p=10^{-3}$ for all quantum operations, which may originate from dominant error sources—whether from measurements or single- or two-qubit physical gates. We sourced the matching widgets from numerical simulations in [40] and simulated a larger $T$-factory, summarizing our findings in table˜1 as an extended $T$-factory Lookup Table.” We recursively search this table to identify a suitable $T$-factory. Notably, table˜1 is part of the user-provided configuration set, $\{\text{config}\}$, and can be customized in RRE.
In $T$-factory Lookup Table”, table˜1, the output probability, $p_{\text{out}}$, determines the logical error rate for the output state by $T$-factories and sets an upper limit on the error rate for any $T$-gate implemented using the auxiliary qubits. The specifications for the number of physical qubits, the dimensions of the 2D patch, and the number of physical cycle sweeps depend on the factory’s construction. Our analysis focused exclusively on selecting $T$-factories that produce $T$-states, thus excluding any entries related to $CCZ$ from the original [40, table 1].
Next, we must include the resource cost of decomposing non-Cliffords into Clifford+$T$ (gate synthesis) to consume graph-state nodes via single-qubit measurements. After completing a preparation sub-schedule, at each sub-step of the consumption schedule, we use one or more purified $T$-resources to perform measurements in the desired basis. The FTQC must move $T$-states from $T$-factories to the auxiliary bus through patch deformations[39, 30] at negligible time cost $O(1)$. We assume that graph nodes either require arbitrary-angle $R_{Z}$-basis measurements or a single $T$-basis measurement. Measurements involving $R_{Z}$-axis rotations are decomposed into sequences of $T$-gates (and Cliffords), with a fixed worst-case length of $L_{\varepsilon}=\lceil c_{0}\log(1/\varepsilon)+c_{1}\rceil$. We determine this length using the gate synthesis methods available to RRE. Here, $\varepsilon$ refers to the diamond-norm error for decompositions into Clifford+TT—see [75, 32] for further details.
Generating each distilled $T$-state in the $T$-factories requires $C$ cycles, and we need one further surface code’s $d$-cycle tock to interact with the graph node and measure it in either the $X$ or $Z$ basis. In this work, we ignore the time cost of any operation that requires $O(1)$ tocks (idling, deforming surface code patches, etc.), assuming we can perform them in parallel with other operations during sub-steps at no additional cost or with negligible delay. Therefore, for each non-Clifford measurement of memory nodes, assuming an $R_{Z}$-basis as an example, the volume increases as follows:
@@
Above, the second term reflects the fact that both graph node consumption and $T$-state distillation occur over a space of $2\tilde{n}_{\text{logical}}+n_{\text{per-leg}}^{\text{modules}}l_{\text{transfer-bus}}$. All nodes with distinct measurement basis types, $T$ and $R_{Z}$, must be measured sequentially.
If we assume the worst-case sequential execution, the upper bound on the full space-time volume becomes:
@@
We define the total number of sequential $T$-states for graph consumption as $N^{\text{seq}}_{\text{consump}}=\left\lceil\frac{n^{T}_{\text{init}}}{n^{\prime}}\right\rceil+L_{\varepsilon}\left\lceil\frac{n^{R_{z}}_{\text{init}}}{n^{\prime}}\right\rceil$. Recall that we assumed a simplified $T$-routing scheme that allows the auxiliary bus to serve up to $n^{\prime}$ simultaneous magic states. FTQC can only consume distinct $T$ and $R_{Z}$ nodes in parallel. The number of sequential $T$-states to distill is $N^{\text{seq}}_{\text{distill}}=\left\lceil\frac{N^{T}_{tot}}{n^{\prime}}\right\rceil$, where the total $T$-state measurements required is defined as $N^{T}_{tot}:=n^{R_{Z}}_{\text{init}}L_{\varepsilon}+n^{T}_{\text{init}}$. The quantity $n^{T}_{tot}$ is the proxy we use for $T$-count and typically becomes larger than $N_{\mathcal{G}}$. We do not include $T$-factory space-time volumes, as their errors subsume into the output probability $p_{out}$ of the distilled state [40, fig. 1]. If the select widget from table˜1 has a second (20-to-4) layer (the second and third widgets), then we must replace $n^{\prime}\rightarrow 4n^{\prime}$ in all space-time equations.
As detailed in section˜II.3.1, a single-cycle step of the surface code (with multiple physical operation layers) has a failure probability $p_{C}$ that characterizes the relationship between the physical and logical error rates, as given in eq.˜1. Consequently, we can express the logical failure probability of a $d$-cycle tock as:
$p_{\text{logical}}=1-(1-p_{C})^{d}.$ (15)
The failure rate for a total volume of these $d$-cycle units, denoted as $V^{\text{sequential}}_{\text{total}}$, must satisfy the following condition:
$p_{\text{algo-fail}}>1-(1-p_{\text{logical}})^{V^{\text{sequential}}_{\text{total}}},$
where $p_{\text{algo-fail}}$ represents the overall algorithm failure rate or error budget for executing the complete circuit $U$. This budget encompasses the entire space-time volume determined by the user within the parameter set $\{\text{config}\}$.
Assuming $p_{C}\ll 1$, we can approximate $(1-p_{C})^{d}\approx 1-p_{C}d$, leading us to conclude that $p_{\text{log-cell}}\approx p_{C}d$. Consequently,
@@
This equation highlights the relationship between the failure rates and the characteristic space-time volume for the complete QEC process.
Subtracting each side of section˜II.4.7 from 1 and taking the logarithm of both sides, we get
@@
If we substitute $V^{\text{sequential}}_{\text{total}}$ from eq.˜14 into eq.˜16, the space-time logarithmic inequality to solve can be expressed as
@@
where $J_{1}=\log(1-p_{\text{algo-fail}})$. The minimum code distance required is the smallest integer $d$ that satisfies eq.˜17. The only free variables are the cycle time of the $T$-factory, i.e. $C$, and $\varepsilon$. We previously extracted $\mathbb{L}_{\text{prep}}$ and $\tilde{n}_{\text{logical}}$ from the Cabaliser and Substrate Scheduler.
@@
To determine $C$ and $\varepsilon$, we perform nested loops to iterate through widgets and then compute $\varepsilon$ of the gate synthesis. The inner loop feeds back into the Clifford+$T$ decomposition, where the failure rate of a $d$-cycle tock, $p_{\text{logical}}$, sets the approximation error as $\varepsilon<p_{\text{logical}}$. By selecting a specific widget, we fix $C$, which also determines the output error rate of the distilled state, $p_{\text{out}}$.
@@
In a compiled quantum algorithm, we aim to ensure that the distilled $T$-states have precisely the same error rates as any other gates in the compiled circuit. A Pauli initialization or measurement occupies a $d$-cycle unit in the space-time volume. Thus, its failure rate is governed by eq.˜15. This error rate must be dominant. Therefore, the error output, $p_{\text{out}}$, from the distillation circuit should be less than the failure rate of this cell. That is, we must ensure that the distillation output is at least as good as a unit cell in the space-time volume. Consequently, we first select the smallest value of $C$ from table˜1 and solve eq.˜17 for $d$. We then verify that, for a physical error rate of $p=10^{-3}$ and our calculated solution for $d$, the following holds for the corresponding $p_{\text{out}}$ entry,
$p_{\text{logical}}=1-(1-p_{C})^{d}>p_{\text{out}}$ (18)
@@
If this condition is unmet, we must choose a better-performing $T$-factory in the outer loop. We then proceed to the next entry on the list and repeat until we identify a $T$-factory that performs at least as well as $p_{\text{logical}}$.
@@
Once a specific $T$-factory is selected, the total number of allocated physical qubits across all modules at any given time and the total intra-modular runtime (including the first subgraph preparation time at step-0, all consumption operations, and delays) can be calculated as follows:
@@
$\begin{split}n&{}_{\text{alloc-phys}}=\\ &4\tilde{n}_{\text{logical}}d^{2}+n^{\text{modules}}_{\text{per-leg}}(n_{\text{T-factories}}Q+2d^{2}l_{\text{transfer-bus}}),\\ t&{}^{\text{total}}_{\text{consump}}=\\ &8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{distill}}+t^{\text{delay}}_{\text{prep}}~.\end{split}$ (19)
@@
In eq.˜19, $t$ denotes the characteristic gate time, which sets the architecture’s quantum tock for all intra-modular operations. The factor of 8 arises because, in our architecture, each surface code sweep for stabilizer syndrome extraction circuits requires at most one initialization, two Hadamards, four entangling gates, and one measurement, all assumed to take time $t$ (see, for instance, [20, fig. 9]). The temporal cost of initializing the first subgraph $g_{0}$ with $L^{0}_{\text{prep}}(S^{0}_{\text{prep}})$ is always nonzero and was explicitly included. FTQC performs the remaining graph preparations and $T$-state distillation operations in parallel with consumption operations. Overall delays may accumulate across time steps, defined as $t^{\text{delay}}_{\text{distill}}=\sum_{i}t^{\text{distill-delay}}_{\text{i}}$ and $t^{\text{delay}}_{\text{prep}}=\sum_{i}t^{\text{prep-delay}}_{\text{i}}$, which we detail below.
@@
To execute each consumption sub-schedule $i$, a specific number of distilled magic states must be present on the module’s $T$-transfer bus. We must introduce a delay if we require more magic states than were accumulated on the $T$-transfer bus during the prior preparation step, i1i-1. The number of $T$ states per module after a preparation step is
@@
$\begin{split}n^{T-\text{per-module}}_{i}=\max\big(\lfloor&\frac{n_{\text{T-factories}}t^{\text{prep}}_{i}}{8tC}\rfloor,l_{\text{transfer-bus}}\big)\\ &\text{for}~i\in\{0,\dots,n_{\text{widgets}}-1\}.\end{split}$ (20)